Two buses leave a station at the same time and travel in opposite directions. One bus travels 8kmh faster than the other. If the two buses are 976 kilometers apart after 4 hours, what is the rate of each bus?
Answer by ikleyn(52778)

.
Two buses leave a station at the same time and travel in opposite directions.
One bus travels 8kmh faster than the other. If the two buses are 976 kilometers apart
after 4 hours, what is the rate of each bus?
~~~~~~~~~~~~~~

Let x be the rate of the slower bus, in kilometers per hour.

Then the rate of the other bus is (x+8) km/h.


Write the total distance equation

    4x + 4*(x+8) = 976.


Simplify and find x

    4x + 4x + 32 = 976

       8x        = 976 - 32 

       8x        = 944

        x        = 944/8 = 118.


ANSWER.  The slower bus speed is  118 km/h.  The faster bus speed is  118+8 = 126 km/h.


CHECK.   4*118 + 4*126 = 976 kilometers, total distance.   ! Precisely correct !

Solved.

---------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.